OpenStudy (anonymous):

Find the radius of convergence and the interval of convergence of the following infinite series: \[\sum_{n=1}^{\infty}{\frac{n!*x^n}{1*3*5*...*(2n-1)}}\]

OpenStudy (anonymous):

:-( I need help with this one...

OpenStudy (unklerhaukus):

are you sure it converges?

OpenStudy (anonymous):

it must (on condition)... it's a power series.

OpenStudy (anonymous):

I got it to be convergent for \(-2\le x \le 2\). You need to test it at the points \(x=-2\) and \(x=2\).

OpenStudy (zarkon):

again...use the ratio test

OpenStudy (anonymous):

\[\left| {a_{n+1} \over a_n} \right|=\left| {n!(n+1)x^{n+1} \over 1 \times 3 \times 5 \times .. \times (2n+1)}. {1 \times 3 \times 5 \times.. (2n-1) \over n!x^n} \right|=\left| {(n+1)x \over 2n+1} \right|.\] \(\left| {n+1 \over 2n+1} \right|\rightarrow{1 \over 2} \text{ as n }\rightarrow \infty\). So it converges for \(-1 < \frac{1}{2}x< 1 \implies -2<x<2\).

OpenStudy (anonymous):

how do I test the bounds of the interval?

OpenStudy (anonymous):

You need now to test the points where \(|\frac{1}{2}x|=1\), that's x=2 and x=-2 using any other test.

OpenStudy (anonymous):

what test would I use? I can't think of any :-(

OpenStudy (anonymous):

Neither can I :D! I'm thinking!

OpenStudy (zarkon):

obvious...use divergence test

OpenStudy (anonymous):

Divergence test would work with \(x=2\), but not with \(x=-2\), right?

OpenStudy (zarkon):

it works with both

OpenStudy (zarkon):

if the limit is not zero or does not converge the series diverges

OpenStudy (anonymous):

what's the divergence test? you mean simply testing the limit of the sequence? it looks inconclusive :-(

OpenStudy (anonymous):

The limit is not \(0\) yeah! I can see that.

OpenStudy (anonymous):

The divergence test states that if the limit of \(a_n\) is not zero, then the series must be divergent.